OpenStudy (anonymous):

Find the determinant of the matrix. -1 2 3 5 0 -2 6 8 1 I got 112. Is that correct?

OpenStudy (anonymous):

=-1( 16) -2(5+12) +3(40)

OpenStudy (anonymous):

=70

OpenStudy (anonymous):

Aw shucks. Thanks.!

OpenStudy (anonymous):

= \[-1\left[\begin{matrix} 0 & -2 \\ 8 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

\[- (2)\left[\begin{matrix}5 & -2\\ 6 & 1\end{matrix}\right] + (3) \left[\begin{matrix}5 & 0 \\ 6 & 8\end{matrix}\right]\]

OpenStudy (anonymous):

where the matricies are determinants

OpenStudy (anonymous):

[0 -1] [5 -2] [5 0] (-1) [8 1] -(2)[6 1] (3)[6 8]

OpenStudy (anonymous):

From Mathematica:\[\text{Det}\left[\left( \begin{array}{ccc} -1 & 2 & 3 \\ 5 & 0 & -2 \\ 6 & 8 & 1 \end{array} \right)\right]=70 \]

OpenStudy (anonymous):

\[\text{Det}\left[\left( \begin{array}{ccc} a & b & c \\ d & e & f \\ g & h & i \end{array} \right)\right]=-c e g+b f g+c d h-a f h-b d i+a e i \]